can you please help me solve this problem: the absolute value of 2a-1 is greater than or equal to 2a+1
Answer by Alan3354(69443)

absolute value of 2a-1 is greater than or equal to 2a+1
|2a-1| >= 2a+1
---------
Case 1:
2a-1 >= 2a+1
-1 >= 1 (Not true, ignore)
------------
Case 2:
-(2a-1) >= 2a+1
-2a + 1 >= 2a+1
1 >= 4a + 1
0 >= 4a
4a <= 0
a <= 0